OpenStudy (anonymous):

Divide 8x^4 – 6x^3 + 7x^2 – 11x + 10 by 2x – 1 using long division. Show all your work. Then explain if 2x – 1 is a factor of the dividend.

OpenStudy (anonymous):

hey @jojo4eva can u help me ?

OpenStudy (whpalmer4):

--------------------------------------------- 2x - 1 | 8x^4-6x^3+7x^2-11x+10 divide the first term of the numerator (\(8x^4)\) by the first term of the denominator \((2x)\) What do you get? That will be the first term of the quotient.

OpenStudy (anonymous):

4x^3

OpenStudy (whpalmer4):

sorry, got called away from the computer. right, \(4x^3\) is the first term of the quotient. Now we multiply that by the denominator (the whole thing), giving us\[4x^3(2x-1) = 8x^4-4x^3\]We subtract that from the numerator: 4x^3 --------------------------------------------- 2x - 1 | 8x^4-6x^3+7x^2-11x+10 - (8x^4 -4x^3) ------------- -2x^3 + 7x^2 - 11x + 10 Now we repeat the process. What is the first term of the remaining numerator \((-2x^3)\) divided by the first term of the denominator \((2x)\)?

OpenStudy (whpalmer4):

\[-x^2(2x-1) = -2x^3+x^2\]We subtract that from the remaining numerator: 4x^3 - x^2 --------------------------------------------- 2x - 1 | 8x^4-6x^3+7x^2-11x+10 - (8x^4 -4x^3) ------------- -2x^3 + 7x^2 - 11x + 10 -(-2x^3 + x^2) --------------- 6x^2 - 11x + 10 we keep doing this procedure until we are left with 0, or a quantity that is not divisible by the denominator. In the latter case, that is the remainder.

OpenStudy (anonymous):

Can you continue @whpalmer4 ?

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

please help yall im extremely bad at this!

OpenStudy (whpalmer4):

Okay, our remaining polynomial is \(6x^2-11x+10\). First term is \(6x^2\). What is \((x^2)/(2x)\)? Again, first term of remaining polynomial divided by first term of denominator.

OpenStudy (anonymous):

1/2 x

OpenStudy (whpalmer4):

\[\frac{6x^2}{2x} = \frac{1}{2}x\]?

OpenStudy (whpalmer4):

sorry, I mistyped a bit, I meant \((6x^2)/(2x)\)

OpenStudy (anonymous):

i did x^2/2x

OpenStudy (anonymous):

3x

OpenStudy (whpalmer4):

Yes, you did. Serves me right for not making you identify the terms to divide :-)

OpenStudy (anonymous):

do i put that above the x term?

OpenStudy (whpalmer4):

Right. Now multiply that result by the denominator, and subtract that result from the remainder of the numerator, what do you get?

OpenStudy (anonymous):

the result of what,

OpenStudy (whpalmer4):

multiply 3x by the denominator. what do you get?

OpenStudy (whpalmer4):

what is the denominator?

OpenStudy (anonymous):

6x^2 -11x+10?

OpenStudy (whpalmer4):

No, the original problem was to divide \(8x^4 – 6x^3 + 7x^2 – 11x + 10 \text{ by } 2x – 1\). Which of those quantities is the denominator (or divisor)?

OpenStudy (anonymous):

2x-1

OpenStudy (anonymous):

so 6x^2 -3x

OpenStudy (whpalmer4):

Correct. So, our latest term of the quotient is \(3x\). Multiply \(3x\) by the denominator to get the latest polynomial to subtract from what is left (the bottom line in the problem so far)

OpenStudy (anonymous):

subtract what i just got from 6x^2 -11x+10?

OpenStudy (whpalmer4):

Yes: 4x^3 - x^2 --------------------------------------------- 2x - 1 | 8x^4-6x^3+7x^2-11x+10 - (8x^4 -4x^3) ------------- -2x^3 + 7x^2 - 11x + 10 -(-2x^3 + x^2) --------------- 6x^2 - 11x + 10

OpenStudy (whpalmer4):

\[6x^2-11x+10 - (6x^2-3x) = \]

OpenStudy (anonymous):

10-8x

OpenStudy (whpalmer4):

Good. that gives us: 4x^3 - x^2 + 3x --------------------------------------------- 2x - 1 | 8x^4-6x^3+7x^2-11x+10 - (8x^4 -4x^3) ------------- -2x^3 + 7x^2 - 11x + 10 -(-2x^3 + x^2) --------------- 6x^2 - 11x + 10 - ( 6x^2 -3x) ------------------- -8x + 10 Okay, divide the first term of the remaining part by the first term of the denominator. What is the result?

OpenStudy (anonymous):

divide -8x by 2x-1?

OpenStudy (whpalmer4):

No, the first term of the remaining part, divided by the first term of the denominator.

OpenStudy (whpalmer4):

Same procedure we've used for the previous 3 terms.

OpenStudy (anonymous):

oh -11x by 2x-1?

OpenStudy (anonymous):

having trouble dividing that

OpenStudy (whpalmer4):

no, divide the first term of the remaining part, which is \(-8x+10\) as you can see from the typed rendition above divide the first term of that by the first term of \(2x-1\)

OpenStudy (anonymous):

i got -4+ 6/2x-1

OpenStudy (whpalmer4):

Yes, so the entire answer is \[4x^3 - x^2 + 3x-4 + \frac{6}{2x-1}\] or \[4x^3 - x^2 + 3x-4 \text{ with remainder } 6\]

OpenStudy (anonymous):

so 2x-1 isn't a factor of the dividend?

OpenStudy (whpalmer4):

nope, a factor will leave a remainder of 0.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks for walking me through that

OpenStudy (anonymous):

i really appreciate it!

OpenStudy (whpalmer4):

That polynomial \[8x^4-6x^3+7x^2-11x+10\]is irreducible, which is a fancy word for "can't be factored"

OpenStudy (anonymous):

okay, thanks

OpenStudy (whpalmer4):

you're welcome! I think (despite the hassle of typing them) these are actually easier to do than long division on numbers, because you never have to worry about your guess for the next term being "too big". You also don't have to worry about getting the digits in the right columns, or any of that. On the other hand, subtracting polynomials with - signs in them is somewhat error-prone for some (myself included).